# MARS: Unleashing the Power of Variance Reduction for Training Large Models

### v2

@@ Abstract
Training deep neural networks—and more recently, large models demands efficient and scalable optimizers. Adaptive gradient algorithms like Adam, AdamW, and their variants have been central to this task. Despite the development of numerous variance reduction algorithms in the past decade aimed at accelerating stochastic optimization in both convex and nonconvex settings, variance reduction has not found widespread success in training deep neural networks or large language models. Consequently, it has remained a less favored approach in modern AI. In this paper, to unleash the power of variance reduction for efficient training of large models, we propose a unified optimization framework, MARS (Make vAriance Reduction Shine), which reconciles preconditioned gradient methods with variance reduction via a scaled stochastic recursive momentum technique. Within our framework, we introduce three instances of MARS that leverage preconditioned gradient updates based on AdamW, Lion, and Shampoo, respectively. We also draw a connection between our algorithms and existing optimizers. Experimental results on training GPT-2 models indicate that MARS consistently outperforms AdamW by a large margin. The implementation of MARS is available at https://github.com/AGI-Arena/MARS.
@@ 1 Introduction
Adaptive gradient methods such as Adam (Kingma and Ba, 2015) and AdamW (Loshchilov and Hutter, 2019) have become the predominant optimization algorithms in deep learning. With the surge of large language models, the majority of the renowned models, including GPT-2 (Radford et al., 2019), GPT-3 (Brown, 2020), PaLM (Chowdhery et al., 2023) and Llama 3 (Dubey et al., 2024) are trained with adaptive gradient methods. Numerous efforts have been made to improve adaptive gradient methods from both first-order and second-order optimization perspectives. For example, You et al. (2019) introduced LAMB, a layerwise adaptation technique that boosts training efficiency for BERT (Devlin, 2018). Using symbolic search, Chen et al. (2023) developed Lion, achieving faster training and reduced memory usage. Liu et al. (2023) designed Sophia, leveraging stochastic diagonal Hessian estimators to accelerate training. Gupta et al. (2018) proposed Shampoo, which performs stochastic optimization over tensor spaces with preconditioning matrices for each dimension. Anil et al. (2020) further refined Shampoo to give a scalable, practical version. Recently, Vyas et al. (2024) showed that Shampoo is equivalent to Adafactor (Shazeer and Stern, 2018) in the eigenbasis of Shampoo’s preconditioner, and introduced SOAP, which stabilizes Shampoo with Adam. Importantly, recent studies (Kaddour et al., 2024; Zhao et al., 2024) have shown that these optimizers perform on par with AdamW in LLM pretraining, yet do not outperform it. This suggests the ongoing challenge in developing adaptive gradient methods superior to Adam and AdamW for large-scale model training.
Since adaptive gradient methods face challenges of high stochastic gradient variance, and language model training inherently involves a high-variance optimization problem (McCandlish et al., 2018), it is natural to consider variance reduction techniques to address this challenge. There exists a large body of literature on variance reduction for stochastic optimization, such as SAG (Roux et al., 2012), SVRG (Johnson and Zhang, 2013), STORM (Cutkosky and Orabona, 2019), which can improve the convergence of stochastic optimization. However, variance reduction has not found widespread success in training deep neural networks or large language models. Defazio and Bottou (2019) discussed why variance reduction can be ineffective in deep learning due to factors such as data augmentation, batch normalization, and dropout, which disrupt the finite-sum structure required by variance reduction principles. Nevertheless, in training language models, data augmentation, batch normalization, and dropout are nowadays rarely used, which opens the door for applying variance reduction techniques in optimizing these models. This naturally leads to the following research question:
@@
In this paper, we answer the above question affirmatively by introducing a novel optimization framework called MARS (Make vAriance Reduction Shine), which incorporates variance reduction into adaptive gradient methods. Notably, we introduce a scaling parameter $\gamma\in(0,1]$ into the stochastic recursive momentum (STORM) (Cutkosky and Orabona, 2019) to adjust the strength of variance reduction and define a new gradient estimator. This gradient estimator undergoes gradient clipping and is subsequently subjected to exponential averaging. When the scaling parameter is set to $\gamma=1$, it recovers the vanilla STORM momentum. In addition, the second-order momentum update is defined by the reweighted intermediate variable. These together ensure optimization stability throughout the training process. We summarize our major contributions of this paper as follows:
@@
  • The MARS framework is versatile, accommodating all existing full matrix or diagonal Hessian approximations. Under this framework, we utilize three distinct designs of the preconditioning matrix, resulting in three specific instances of our MARS framework: MARS-AdamW, MARS-Lion, and MARS-Shampoo. Each variant demonstrates compatibility with their corresponding preconditioning in AdamW, Lion, and Shampoo, showing that MARS can seamlessly integrate with and do variance-reduction on these established methods.
  • Theoretically, we prove that MARS achieves an $\mathcal{O}(T^{-2/3})$ convergence rate to the first-order stationary points, which improves upon the $\mathcal{O}(T^{-1/2})$ convergence rate of Adam/AdamW  (Zhou et al., 2018; Chen et al., 2018b; Huang et al., 2021). Here $T$ is the number of iterations. Moreover, we precisely characterize the dependence of the upper bound on the scaling parameter $\gamma$, which controls the strength of variance reduction. When $\gamma=1$, the proposed momentum reduces to the STORM momentum, and the convergence rate recovers that of SuperAdam (Huang et al., 2021). Notably, an optimal $\gamma$ schedule exists that enables a faster convergence rate, demonstrating that the flexibility in choosing $\gamma$ leads to a tighter upper bound.
  • Empirically, we evaluated MARS on GPT-2 fine-tuning tasks using the OpenWebText dataset. It demonstrates superior performance on GPT-2 large: AdamW requires 50 billion tokens to reaches a validation loss of 2.58, whereas MARS only requires 28 billion tokens, and it achieves a final validation loss of 2.51. Furthermore, on the downstream task Hellaswag, MARS improved accuracy to 44.64%, outperforming AdamW’s 41.70% after training on 50 billion tokens.
Notations
In this paper, we assume $\mathbf{x}_{t}$ denotes the parameter of the language model at step $t$ and $\bm{\xi}_{1},...,\bm{\xi}_{T}\in\mathbf{\Xi}$ are a sequence of independent random variables which denote the training data for each step. For some objective function $f$ that is differentiable, we assume $\mathbb{E}[f(\mathbf{x},\bm{\xi}_{t})|\mathbf{x}]=F(\mathbf{x})$ for $\forall\mathbf{x},~{}\forall t$. In our algorithm, the training data of the current step $\bm{\xi}_{t}$ and previous step $\bm{\xi}_{t-1}$ are used for attaining different gradient for the same parameter $\mathbf{x}_{t}$, so we just explicitly indicate these variables for function $f$.
@@ 2 Preliminaries
where $\eta_{t}>0$ is the learning rate or step size. SGD needs $\mathcal{O}(\varepsilon^{-4})$ stochastic gradient evaluations (i.e., gradient complexity or incremental first-order oracle complexity) to find a $\epsilon$-approximate first-order stationary points, i.e., $\|\nabla F(\mathbf{x})\|_{2}\leq\epsilon$ (Ghadimi and Lan, 2013).
To accelerate the convergence of SGD, variance reduction techniques have been extensively researched in both the machine learning and optimization communities over the past decade, resulting in numerous algorithms for convex optimization—such as SAG (Roux et al., 2012), SVRG (Johnson and Zhang, 2013), SAGA (Defazio et al., 2014), and SARAH (Nguyen et al., 2017a)—as well as for nonconvex optimization, including SVRG (Allen-Zhu and Yuan, 2016; Reddi et al., 2016), SNVRG (Zhou et al., 2020), SPIDER (Fang et al., 2018), and STORM (Cutkosky and Orabona, 2019), among others. Notably, for nonconvex optimization, SNVRG (Zhou et al., 2020), SPIDER (Fang et al., 2018) and STORM (Cutkosky and Orabona, 2019) can improve the gradient complexity of SGD from $\mathcal{O}(\varepsilon^{-4})$ to $\mathcal{O}(\varepsilon^{-3})$, demonstrating a provable advantage.
At the heart of variance reduction techniques is a variance-reduced stochastic gradient, exemplified by the method proposed by Johnson and Zhang (2013) as follows:
@@
Subsequently, Cutkosky and Orabona (2019) introduced Stochastic Recursive Momentum (STORM), a variant of standard momentum with an additional term, achieving the same convergence rate as SPIDER while eliminating the need for periodic anchoring:
$\mathbf{m}_{t}$ $=\beta_{1}\mathbf{m}_{t-1}+(1-\beta_{1})\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})+\beta_{1}\big{(}\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})-\nabla f(\mathbf{x}_{t-1},\bm{\xi}_{t})\big{)}$ (2.4)
@@
$\mathbf{m}_{t}$ $=\beta_{1}\mathbf{m}_{t-1}+(1-\beta_{1})\Big{[}\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})+\underbrace{\frac{\beta_{1}}{1-\beta_{1}}\big{(}\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})-\nabla f(\mathbf{x}_{t-1},\bm{\xi}_{t})\big{)}}_{\text{gradient correction}}\Big{]}.$ (2.5)
@@ 3 Method
In this section, we introduce MARS (Make vAriance Reduction Shine), a family of precondtioned optimization algorithms that perform variance reduction in gradient estimation.
@@ 3.1 MARS Framework
Preconditioned Variance Reduction.
Variance reduction methods achieve faster convergence than SGD, yet identifying optimal learning rates remains a practical challenge. Particularly, different parameters often exhibit varying curvatures, requiring tailored learning rates for each. One approach to addressing this issue is to use the Hessian matrix to precondition gradient updates, integrating curvature information into the updates. The idea stems from minimizing the second-order Taylor expansion at $\mathbf{x}_{t}$:
$F(\mathbf{x}_{t+1})$ $\approx F(\mathbf{x}_{t})+\nabla F(\mathbf{x}_{t})(\mathbf{x}_{t+1}-\mathbf{x}_{t})+\frac{1}{2}(\mathbf{x}_{t+1}-\mathbf{x}_{t})^{\top}\nabla^{2}F(\mathbf{x}_{t})(\mathbf{x}_{t+1}-\mathbf{x}_{t}),$ (3.1)
@@
$\mathbf{m}_{t}$ $=\beta_{1}\mathbf{m}_{t-1}+(1-\beta_{1})\Big{[}\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})+\frac{\beta_{1}}{1-\beta_{1}}\big{(}\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})-\nabla f(\mathbf{x}_{t-1},\bm{\xi}_{t})\big{)}\Big{]},$ (3.3)
@@
SuperAdam (Huang et al., 2021) also incorporates the STORM into the design of adaptive gradient methods. However, their precondition matrix can be viewed as a special case of our general framework. SuperAdam’s design focuses on diagonal precondition matrix and draws heavily from the design used in Adam (Kingma and Ba, 2015), AdaGrad-Norm (Ward et al., 2020), and AdaBelief (Zhuang et al., 2020). Furthermore, their preconditioner matrix is designed following Adam’s structure but does not account for the revised definition of variance-reduced momentum, resulting in a significant mismatch between the first-order and second-order momentum. We will further clarify these differences when discussing specific instances of our framework.
Algorithm Design.
In practice, alongside our preconditioned variance-reduced update (3.3), we introduce a scaling parameter $\gamma_{t}$ to control the scale of gradient correction in variance reduction. We also introduce a new gradient estimator $\mathbf{c}_{t}$, which is the combination of stochastic gradient and the scaled gradient correction term:
$\mathbf{c}_{t}=\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})+\underbrace{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\gamma_{t}}\frac{\beta_{1}}{1-\beta_{1}}\big{(}\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})-\nabla f(\mathbf{x}_{t-1},\bm{\xi}_{t})\big{)}}_{\text{scaled gradient correction}}.$
@@
Why $\gamma_{t}$ improves convergence
A similar idea of adjusting variance reduction strength has been proposed in Yin et al. (2023) for SVRG. For the STORM updates used in our work, we provide an intuition by the estimation of
$\mathbb{E}[X+(\gamma Y-\mathbb{E}Y)]^{2},$
where $X$ and $Y$ are two independent random variables. [^1] In our case, $Y=\nabla f(\mathbf{x}_{t+1},\bm{\xi}_{t+1})-\nabla f(\mathbf{x}_{t},\bm{\xi}_{t+1})$. See Proof of Lemma B.2 for details.
  • When $\gamma=1$, this simplifies to $\mathbb{E}[X+(Y-\mathbb{E}Y)]^{2}=\mathbb{E}X^{2}+\mathrm{Var}(Y)$.
  • When $\gamma=0$ it reduces to $\mathbb{E}[X-\mathbb{E}Y]^{2}$, which can be significantly larger than choosing $\gamma=1$ when $\mathbb{E}XY<0$.
  • There exists an optimal choice of $\gamma^{*}$, given by $\gamma^{*}=1-\frac{\mathbb{E}XY+\mathrm{Var}(Y)}{\mathbb{E}Y^{2}}$, such that
    $\mathbb{E}[X+(\gamma Y-\mathbb{E}Y)]^{2}$ $=\mathbb{E}[X+(Y-\mathbb{E}Y)]^{2}-\big{(}\mathbb{E}XY+\mathrm{Var}(Y)\big{)}^{2}/\mathbb{E}Y^{2},$
@@
which yields a smaller variance compared to any other choice of $\gamma$. We provide a convergence analysis of Algorithm 1 in Theorem 4.4 in Section 4. We prove that under standard assumptions, MARS achieves a superior convergence rate of $\mathcal{O}(T^{-2/3})$, outperforming the $\mathcal{O}(T^{-1/2})$ rate attainable by AdamW and its variants (Zhou et al., 2018; Chen et al., 2018b; Huang et al., 2021).
Full Matrix Approximation.
@@
In practice, calculating the Hessian matrix is computationally expensive or even intractable due to the complexity of second-order differentiation and the significant memory cost of storing $\mathbf{H}_{t}$, especially when the parameters in a neural network constitute a high-dimensional matrix. Many existing algorithms employ various approximations of the Hessian. For instance, K-FAC (Martens and Grosse, 2015) and Shampoo (Gupta et al., 2018) approximate the Gauss-Newton component of the Hessian (also known as the Fisher information matrix), using a layerwise Kronecker product approximation (Morwani et al., 2024). Additionally, Sophia (Liu et al., 2023) suggests using Hutchinson’s estimator or the Gauss-Newton-Barlett estimator for approximating the Hessian. We take various designs of the preconditioning matrix into account and broaden the definition of $\mathbf{H}_{t}$ in (3.4) to encompass various specifically designed preconditioning matrix in the rest of the paper.
Diagonal Matrix Approximation.
@@
Even when using approximated Hessian matrices, second-order algorithms mentioned above remain more computationally intensive compared to first-order gradient updates. Thus, another line of research focuses on approximating the Hessian matrix through diagonal matrices, as seen in optimization algorithms like AdaGrad (Duchi et al., 2011), RMSProp (Tieleman, 2012), AdaDelta (Zeiler, 2012), Adam (Kingma and Ba, 2015) and AdamW (Loshchilov and Hutter, 2019), etc. This approach to diagonal preconditioning effectively transforms the updates into a first-order method, assigning adaptive learning rates to each gradient coordinate. For example, in AdaGrad (Duchi et al., 2011), the preconditioned matrix is defined by:

### v3

@@ Abstract
Training deep neural networks—and more recently, large models demands efficient and scalable optimizers. Adaptive gradient algorithms like Adam, AdamW, and their variants have been central to this task. Despite the development of numerous variance reduction algorithms in the past decade aimed at accelerating stochastic optimization in both convex and nonconvex settings, variance reduction has not found widespread success in training deep neural networks or large language models. Consequently, it has remained a less favored approach in modern AI. In this paper, to unleash the power of variance reduction for efficient training of large models, we propose a unified optimization framework, MARS (Make vAriance Reduction Shine), which reconciles preconditioned gradient methods with variance reduction via a scaled stochastic recursive momentum technique. Within our framework, we introduce three instances of MARS that leverage preconditioned gradient updates based on AdamW, Lion, and Shampoo, respectively. We also draw a connection between our algorithms and existing optimizers. Experimental results on training GPT-2 models indicate that MARS consistently outperforms AdamW by a large margin.
Optimization, Large language model
@@ 1 Introduction
Adaptive gradient methods such as Adam (Kingma & Ba, 2015) and AdamW (Loshchilov & Hutter, 2019) have become the predominant optimization algorithms in deep learning. With the surge of large language models, the majority of the renowned models, including GPT-2 (Radford et al., 2019), GPT-3 (Brown, 2020), PaLM (Chowdhery et al., 2023) and Llama 3 (Dubey et al., 2024) are trained with adaptive gradient methods. Numerous efforts have been made to improve adaptive gradient methods from both first-order and second-order optimization perspectives. For example, You et al. (2019) introduced LAMB, a layerwise adaptation technique that boosts training efficiency for BERT (Devlin, 2018). Using symbolic search, Chen et al. (2023) developed Lion, achieving faster training and reduced memory usage. Liu et al. (2023) designed Sophia, leveraging stochastic diagonal Hessian estimators to accelerate training. Gupta et al. (2018) proposed Shampoo, which performs stochastic optimization over tensor spaces with preconditioning matrices for each dimension. Anil et al. (2020) further refined Shampoo to give a scalable, practical version. Recently, Vyas et al. (2024) showed that Shampoo is equivalent to Adafactor (Shazeer & Stern, 2018) in the eigenbasis of Shampoo’s preconditioner, and introduced SOAP, which stabilizes Shampoo with Adam. Importantly, recent studies (Kaddour et al., 2024; Zhao et al., 2024) have shown that these optimizers perform on par with AdamW in LLM pretraining, yet do not outperform it. This suggests the ongoing challenge in developing adaptive gradient methods superior to Adam and AdamW for large-scale model training.
Since adaptive gradient methods face challenges of high stochastic gradient variance, and language model training inherently involves a high-variance optimization problem (McCandlish et al., 2018), it is natural to consider variance reduction techniques to address this challenge. There exists a large body of literature on variance reduction for stochastic optimization, such as SAG (Roux et al., 2012), SVRG (Johnson & Zhang, 2013), STORM (Cutkosky & Orabona, 2019), which can improve the convergence of stochastic optimization. However, variance reduction has not found widespread success in training deep neural networks or large language models. Defazio & Bottou (2019) discussed why variance reduction can be ineffective in deep learning due to factors such as data augmentation, batch normalization, and dropout, which disrupt the finite-sum structure required by variance reduction principles. Nevertheless, in training language models, data augmentation, batch normalization, and dropout are nowadays rarely used, which opens the door for applying variance reduction techniques in optimizing these models. This naturally leads to the following research question:
@@
In this paper, we answer the above question affirmatively by introducing a novel optimization framework called MARS (Make vAriance Reduction Shine), which incorporates variance reduction into adaptive gradient methods. Notably, we introduce a scaling parameter into the stochastic recursive momentum (STORM) (Cutkosky & Orabona, 2019) to adjust the strength of variance reduction and define a new gradient estimator. This gradient estimator undergoes gradient clipping and is subsequently subjected to exponential averaging. When the variance reduction strength is set to $1$, it recovers the vanilla STORM momentum. In addition, the second-order momentum update is defined by the reweighted intermediate variable. These together ensure optimization stability throughout the training process. We summarize our major contributions of this paper as follows:
@@
  • The MARS framework is versatile, accommodating all existing full matrix or diagonal Hessian approximations. Under this framework, we utilize three distinct designs of the preconditioning matrix, resulting in three specific instances of our MARS framework: MARS-AdamW, MARS-Lion, and MARS-Shampoo. Each variant demonstrates compatibility with their corresponding preconditioning in AdamW, Lion, and Shampoo, showing that MARS can seamlessly integrate with and do variance reduction on these established methods.
  • Empirically, we evaluated MARS on GPT-2 fine-tuning tasks using the OpenWebText dataset. It demonstrates superior performance on GPT-2 large: AdamW requires 50 billion tokens to reach a validation loss of 2.58, whereas MARS only requires 28 billion tokens, and it achieves a final validation loss of 2.51. Furthermore, on the downstream task Hellaswag, MARS improved accuracy to 44.64%, outperforming AdamW’s 41.70% after training on 50 billion tokens. And the code is available at https://github.com/AGI-Arena/MARS.
Notations In this paper, we assume $\mathbf{x}_{t}$ denotes the parameter of the language model at step $t$ and $\bm{\xi}_{1},...,\bm{\xi}_{T}\in\mathbf{\Xi}$ are a sequence of independent random variables which denote the training data for each step. For some objective function $f$ that is differentiable, we assume $\mathbb{E}[f(\mathbf{x},\bm{\xi}_{t})|\mathbf{x}]=F(\mathbf{x})$ for $\forall\mathbf{x},~{}\forall t$. In our algorithm, the training data of the current step $\bm{\xi}_{t}$ and previous step $\bm{\xi}_{t-1}$ are used for attaining different gradient for the same parameter $\mathbf{x}_{t}$, so we just explicitly indicate these variables for function $f$.
@@ 2 Preliminaries
where $\eta_{t}>0$ is the learning rate or step size. SGD needs $\mathcal{O}(\varepsilon^{-4})$ stochastic gradient evaluations (i.e., gradient complexity or incremental first-order oracle complexity) to find a $\epsilon$-approximate first-order stationary points, i.e., $\|\nabla F(\mathbf{x})\|_{2}\leq\epsilon$ (Ghadimi & Lan, 2013).
To accelerate the convergence of SGD, variance reduction techniques have been extensively researched in both the machine learning and optimization communities over the past decade, resulting in numerous algorithms for convex optimization—such as SAG (Roux et al., 2012), SVRG (Johnson & Zhang, 2013), SAGA (Defazio et al., 2014), and SARAH (Nguyen et al., 2017a)—as well as for nonconvex optimization, including SVRG (Allen-Zhu & Yuan, 2016; Reddi et al., 2016), SNVRG (Zhou et al., 2020), SPIDER (Fang et al., 2018), and STORM (Cutkosky & Orabona, 2019), among others. Notably, for nonconvex optimization, SNVRG (Zhou et al., 2020), SPIDER (Fang et al., 2018) and STORM (Cutkosky & Orabona, 2019) can improve the gradient complexity of SGD from $\mathcal{O}(\varepsilon^{-4})$ to $\mathcal{O}(\varepsilon^{-3})$, demonstrating a provable advantage.
At the heart of variance reduction techniques is a variance-reduced stochastic gradient, exemplified by the method proposed by Johnson & Zhang (2013) as follows:
@@
Subsequently, Cutkosky & Orabona (2019) introduced Stochastic Recursive Momentum (STORM), a variant of standard momentum with an additional term, achieving the same convergence rate as SPIDER while eliminating the need for periodic anchoring:
$\mathbf{m}_{t}$ $=\beta_{1}\mathbf{m}_{t-1}+(1-\beta_{1})\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})$
$+\beta_{1}\big{(}\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})-\nabla f(\mathbf{x}_{t-1},\bm{\xi}_{t})\big{)}$ (2.4)
@@
$\mathbf{m}_{t}$ $=\beta_{1}\mathbf{m}_{t-1}+(1-\beta_{1})\Big{[}\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})$
$+\underbrace{\frac{\beta_{1}}{1-\beta_{1}}\big{(}\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})-\nabla f(\mathbf{x}_{t-1},\bm{\xi}_{t})\big{)}}_{\text{gradient correction}}\Big{]}.$ (2.5)
@@ 3 Method
In this section, we introduce MARS (Make vAriance Reduction Shine), a family of preconditioned optimization algorithms that perform variance reduction in gradient estimation.
@@ 3.1 MARS Framework
Preconditioned Variance Reduction. Variance reduction methods achieve faster convergence than SGD, yet identifying optimal learning rates remains a practical challenge. Particularly, different parameters often exhibit varying curvatures, requiring tailored learning rates for each. One approach to addressing this issue is to use the Hessian matrix to precondition gradient updates, integrating curvature information into the updates. The idea stems from minimizing the second-order Taylor expansion at $\mathbf{x}_{t}$:
$F(\mathbf{x}_{t+1})$ $\approx F(\mathbf{x}_{t})+\nabla F(\mathbf{x}_{t})(\mathbf{x}_{t+1}-\mathbf{x}_{t})$
$+\frac{1}{2}(\mathbf{x}_{t+1}-\mathbf{x}_{t})^{\top}\nabla^{2}F(\mathbf{x}_{t})(\mathbf{x}_{t+1}-\mathbf{x}_{t}),$ (3.1)
@@
$\mathbf{m}_{t}$ $=\beta_{1}\mathbf{m}_{t-1}+(1-\beta_{1})\Big{[}\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})$
$+\frac{\beta_{1}}{1-\beta_{1}}\big{(}\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})-\nabla f(\mathbf{x}_{t-1},\bm{\xi}_{t})\big{)}\Big{]},$ (3.3)
@@
SuperAdam (Huang et al., 2021) also incorporates the STORM into the design of adaptive gradient methods. However, their precondition matrix can be viewed as a special case of our general framework. SuperAdam’s design focuses on diagonal precondition matrix and draws heavily from the design used in Adam (Kingma & Ba, 2015), AdaGrad-Norm (Ward et al., 2020), and AdaBelief (Zhuang et al., 2020). Furthermore, their preconditioner matrix is designed following Adam’s structure but does not account for the revised definition of variance-reduced momentum, resulting in a significant mismatch between the first-order and second-order momentum. We will further clarify these differences when discussing specific instances of our framework.
Algorithm Design. In practice, alongside our preconditioned variance-reduced update (3.3), we introduce a scaling parameter $\gamma_{t}$ to control the scale of gradient correction in variance reduction. We also introduce a new gradient estimator $\mathbf{c}_{t}$, which is the combination of stochastic gradient and the scaled gradient correction term:
$\mathbf{c}_{t}=\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})+\underbrace{{\color[rgb]{1,0,0}\gamma_{t}}\frac{\beta_{1}}{1-\beta_{1}}\big{(}\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})-\nabla f(\mathbf{x}_{t-1},\bm{\xi}_{t})\big{)}}_{\text{scaled gradient correction}}.$
@@
Why $\gamma_{t}$ improves convergence A similar idea of adjusting the strength of variance reduction has been proposed by Yin et al. (2023) in the context of SVRG. This approach originates from a classical line of work on control variates (Asmussen & Glynn, 2007; Lavenberg et al., 1977).
In the standard control variates setting, one considers the estimator:
$\mathbb{E}\left[X-\mathbb{E}[X]-{\color[rgb]{1,0,0}\gamma}(Y-\mathbb{E}[Y])\right]^{2}$
$=\mathrm{Var}(X)-2{\color[rgb]{1,0,0}\gamma}\,\mathbb{E}[(X-\mathbb{E}[X])(Y-\mathbb{E}[Y])]+{\color[rgb]{1,0,0}\gamma}^{2}\,\mathrm{Var}(Y),$
which admits an optimal choice of ${\color[rgb]{1,0,0}\gamma}$ that minimizes the variance:
$\arg\min_{{\color[rgb]{1,0,0}\gamma}}\mathbb{E}\left[X-\mathbb{E}[X]-{\color[rgb]{1,0,0}\gamma}(Y-\mathbb{E}[Y])\right]^{2}$
$=\frac{\mathbb{E}[(X-\mathbb{E}[X])(Y-\mathbb{E}[Y])]}{\mathrm{Var}(Y)}.$
In the context of STORM updates used in our work, the update rule includes both stochastic gradients and recursive momentum terms. Let us define:
$X$ $=(1-\beta)\nabla f(\mathbf{x}_{t+1},\bm{\xi}_{t+1})$
$Y$ $=\beta\left[\nabla f(\mathbf{x}_{t+1},\bm{\xi}_{t+1})-\nabla f(\mathbf{x}_{t},\bm{\xi}_{t+1})\right]$
$Z_{t}$ $=\mathbf{m}_{t}-\nabla F(\mathbf{x}_{t})$
and let $U:=X-\mathbb{E}[X]+Z_{t}$. The updated $Z_{t+1}$ at step $t+1$ is of the form $U+({\color[rgb]{1,0,0}\gamma}Y-\mathbb{E}[Y])$, whose squared expectation we aim to minimize.
The optimal choice of ${\color[rgb]{1,0,0}\gamma}$ in this setting is:
${\color[rgb]{1,0,0}\gamma^{*}}$ $=1-\frac{\mathbb{E}[UY]+\mathrm{Var}(Y)}{\mathbb{E}[Y^{2}]}.$
With this choice, we obtain a variance reduction:
$\mathbb{E}\left[U+({\color[rgb]{1,0,0}\gamma^{*}}Y-\mathbb{E}[Y])\right]^{2}$ $=\mathbb{E}\left[U+(Y-\mathbb{E}[Y])\right]^{2}$
$\quad-\frac{\left(\mathbb{E}[UY]+\mathrm{Var}(Y)\right)^{2}}{\mathbb{E}[Y^{2}]},$
which is strictly smaller than the variance under any non-optimal choice of $\gamma$. Thus, dynamically tuning $\gamma_{t}$ improves the variance of the gradient estimator, leading to better convergence behavior. A full convergence analysis is provided in Appendix B.2.
@@
We provide the convergence analysis of Algorithm 1 in Theorem B.5 in Appendix B.2. We prove that under standard assumptions, MARS achieves a superior convergence rate of $\mathcal{O}(T^{-1/3})$, outperforming the $\mathcal{O}(T^{-1/4})$ rate attainable by AdamW.
@@
Full Matrix Approximation. In practice, calculating the Hessian matrix is computationally expensive or even intractable due to the complexity of second-order differentiation and the significant memory cost of storing $\mathbf{H}_{t}$, especially when the parameters in a neural network constitute a high-dimensional matrix. Many existing algorithms employ various approximations of the Hessian. For instance, K-FAC (Martens & Grosse, 2015) and Shampoo (Gupta et al., 2018) approximate the Gauss-Newton component of the Hessian (also known as the Fisher information matrix), using a layerwise Kronecker product approximation (Morwani et al., 2024). Additionally, Sophia (Liu et al., 2023) suggests using Hutchinson’s estimator or the Gauss-Newton-Barlett estimator for approximating the Hessian. We take various designs of the preconditioning matrix into account and broaden the definition of $\mathbf{H}_{t}$ in (3.4) to encompass various specifically designed preconditioning matrix in the rest of the paper.
@@
Diagonal Matrix Approximation. Even when using approximated Hessian matrices, second-order algorithms mentioned above remain more computationally intensive compared to first-order gradient updates. Thus, another line of research focuses on approximating the Hessian matrix through diagonal matrices, as seen in optimization algorithms like AdaGrad (Duchi et al., 2011), RMSProp (Tieleman, 2012), AdaDelta (Zeiler, 2012), Adam (Kingma & Ba, 2015) and AdamW (Loshchilov & Hutter, 2019), etc. This approach to diagonal preconditioning effectively transforms the updates into a first-order method, assigning adaptive learning rates to each gradient coordinate. For example, in AdaGrad (Duchi et al., 2011), the preconditioned matrix is defined by: